OpenStudy (anonymous):

Factor 3x2 - 6x - 9 completely. 3(x - 1)(x + 3) 3(x + 1)(x - 3) (3x + 1)(x - 3) (3x +1)(x - 3)

OpenStudy (anonymous):

Compare it with : \(ax^2 + bx + c\) You will have : a = 3, b = -6 and c = -9

OpenStudy (anonymous):

Always firstly you multiply a with c.. Can you tell me what will you get for \(a \times c\) = ??

OpenStudy (anonymous):

I need your cooperation @andy11

OpenStudy (anonymous):

sorry i had to do somthing. it would = -27

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

So, here is the actual task now.. You have to find factors of -27, so that you get b ie -6..

OpenStudy (anonymous):

Tell me what will be the two numbers if you multiply them you get -27 and if you add them you get -6??

OpenStudy (anonymous):

Getting my point??

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yeah im getting it sorry im just a little slow

OpenStudy (anonymous):

Take your time, I can surely wait..

OpenStudy (anonymous):

You have to choose two numbers..

OpenStudy (anonymous):

Like I choose two numbers as : 1 and -27.. Here, multiplication will be give my -27 but the sum is not equal to -6 as : \(1+( - 27) = -26\)